### 28 Apr 2026 v1

# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

Abstract
Data assimilation (DA) integrates observational information with model predictions to improve state estimation in complex systems. While filtering provides the basis for online forecasts by using only past and present observations, it can exhibit delays and biases when the underlying dynamics evolve rapidly or undergo regime transitions. Smoothing, which additionally incorporates future observations, provides a natural pipeline for hindcasting and reanalysis that yields an uncertainty reduction beyond the filter. This paper introduces an ensemble Kalman–Bucy smoother (EnKBS) for continuous-time DA of nonlinear dynamical systems, where the smoother’s conditional distributions are reconstructed using ensemble moments. The result is a derivative-free framework that does not require explicit computation of tangent-linear or adjoint models, which converges to the exact smoother solution at the infinite-ensemble limit for a wide class of complex systems. Incorporating standard regularization techniques for high-dimensional systems, such as covariance localization and inflation, the skill of the EnKBS is demonstrated in various important scientific problems. By integrating future observations, which reveal the underlying causal mechanisms for retrospective state updates, the EnKBS is used for Bayesian-based inference of causal relationships and their temporal influence range in a dyadic trigger-feedback model and the development of a causality-driven iterative learning algorithm that identifies the structure and recovers the hidden parameters of a nonlinear reduced-order model mimicking midlatitude atmospheric circulation. Notably, both tasks remain effective with an ensemble size of $O(10)$ under partial observations, suggesting that EnKBS can support the instantaneous discovery of high-dimensional complex systems over time.
keywords:
Continuous-time data assimilation , Ensemble Kalman–Bucy filter and smoother , Parameter estimation , Causal inference , Model identification
[…]
1 Introduction
Complex dynamical systems arise in many scientific areas, such as atmospheric and ocean science, fluid dynamics, geophysics, and neuroscience [1, 2, 3, 4, 5]. Such systems are characterized by high dimensionality, strong nonlinearity, and nontrivial driving mechanisms [3, 6, 7, 8].
A central mathematical problem when studying complex dynamical systems is reconstructing the full state of the system from partial and noisy observations, which is essential for system reanalysis, uncertainty quantification, forecasting, and optimal control [9, 10, 11, 12, 13]. More importantly, accurate state estimation is often an essential prerequisite for model identification and the discovery of the underlying causal structure [14, 15, 16].
However, accurate state estimation is challenging. Uncertainties arise primarily from two sources: measurement noise due to imperfect observation tools or methodologies, and model error stemming from incomplete physical understanding or limited computational resolution. In chaotic systems, these uncertainties are amplified over time through multiscale nonlinear interactions, so state estimates that rely entirely on model simulations can quickly become unreliable [11, 17, 18]. In contrast, purely data-driven approaches that rely on statistical interpolation of time-series data often fail to incorporate temporal dynamical knowledge and to enforce physics constraints [15, 19]. A strategy is therefore needed that integrates observational information with model dynamics while addressing the limitations of each. Data assimilation (DA) has been the de facto approach, leading to a powerful framework that systematically integrates observations with dynamics to improve state estimation and reduce uncertainty and biases [20, 21, 12]. Over the past few decades, DA and its variants have been extensively developed and deployed across various disciplines, with notable successes in numerical weather prediction and medical imaging [22, 23, 24].
Broadly speaking, classical DA algorithms are categorized into filters and smoothers. Filters estimate the state using observations up to and including the present, whereas smoothers incorporate information from the entire observation window, including past, present, and future observations. While filtering is the standard approach for real-time forecasting, smoothing is theoretically more accurate for retrospective analysis [25, 26, 27]. On the one hand, smoothing systematically improves estimation error and reduces uncertainty [25, 28]. On the other hand, smoothing can mitigate the detection delay that filters often exhibit at the onset of extreme events. This distinction becomes particularly important in systems where the underlying dynamics are driven by hidden forcing mechanisms. As a result, filters tend to detect the driving signal only after its effects have been observed [29]. In contrast, by leveraging future information, smoothers can more accurately identify the onset and persistence of the underlying driving mechanisms [16, 30, 31].
Although a wide range of DA algorithms has been developed, many methods face theoretical and practical limitations. The Kalman filter and the Rauch–Tung–Striebel (RTS) smoother provide optimal solutions, but only for linear-Gaussian systems [32, 33]. In complex dynamical systems, however, strongly non-Gaussian features such as skewness and heavy tails often violate the Gaussian assumptions underlying the classical theory [34]. For general nonlinear dynamics, variational DA approaches typically provide deterministic state estimates that do not quantify uncertainty, while particle methods suffer from severe computational burdens in high dimensions due to the curse of dimensionality [35, 36, 37]. Consequently, ensemble Kalman methods have emerged as an effective alternative [38, 39]. By approximating system statistics with a finite ensemble, these methods provide state estimates that explicitly quantify uncertainty while remaining derivative-free and computationally feasible in many high-dimensional settings. In practice, the ensemble size (typically $O(10^{2}))$ is often chosen independently of the system’s effective dimension [40, 41]. Due to these advantages, ensemble methods have been widely adopted in large-scale operational applications, including numerical weather prediction, ocean science, geophysics, and fluid dynamics [42, 43, 40, 44]. Most existing ensemble filtering and smoothing algorithms, however, are formulated in discrete time, reflecting the sequential structure of forecast-analysis cycles. Efficient continuous-time smoothers for nonlinear systems remain comparatively less explored, despite their importance for systems naturally described by stochastic differential equations. Existing formulations, such as the continuous-time unscented RTS smoother [45], directly propagate the smoother mean and covariance. In contrast, our ensemble-based smoother evolves stochastic trajectories, providing conditional samples for downstream tasks such as model discovery, while naturally admitting covariance localization for high-dimensional systems.
In this paper, we develop a continuous-time ensemble Kalman–Bucy smoother (EnKBS) for efficient state estimation and trajectory sampling in complex nonlinear dynamical systems. To address high-dimensional settings, the method incorporates standard regularization techniques, including covariance localization and inflation, which mitigate sampling errors, improve numerical stability, and prevent ensemble divergence in nonlinear models [46, 47, 48]. Beyond state estimation, we demonstrate that EnKBS provides a powerful tool for dynamical discovery. We incorporate EnKBS into a Bayesian causal inference framework that identifies causal relationships by tracing driving signals backwards from their observed effects [16]. Numerical tests are presented using a nonlinear dyad model exhibiting extreme events, for which closed-form filtering and smoothing distributions are available [16, 49]. We verify asymptotic convergence of the method to the exact analytical solution and show that it remains accurate and statistically consistent even with an ensemble size of $O(10)$. We further integrate EnKBS into a causality-based iterative learning algorithm for partially observed systems, where EnKBS-based trajectory sampling enables the recovery of hidden parameters and identification of the underlying model structure [15].
The rest of this paper is organized as follows. Section 2 formulates the problem and introduces the proposed continuous-time EnKBS, together with regularization strategies for high-dimensional systems. Section 3 evaluates EnKBS on the high-dimensional Lorenz-96 system [18] and presents applications to assimilative causal inference (ACI) and causality-based model discovery [15, 16]. We conclude the paper with a discussion in Section 4. Mathematical derivations of the EnKBS and its connection to the discrete-time RTS smoother in a formal time limit are deferred to A. Finally, B shows that the proposed EnKBS recovers the exact mean and covariance equations of the classical continuous-time RTS smoother in the linear-Gaussian setting.
2 The Continuous-Time Ensemble Kalman–Bucy Smoother
2.1 Problem Setup
We consider a general class of continuous-time nonlinear systems. Let $\mathbf{x}(t)\in\mathbb{R}^{n_{\mathbf{x}}}$ denote the hidden states and $\mathbf{y}(t)\in\mathbb{R}^{n_{\mathbf{y}}}$ denote the observations (data). The dynamics are governed by the following stochastic differential equations (SDEs) [34, 50]:
$\,\mathrm{d}\mathbf{x}$ $=\mathbf{f}(\mathbf{x},\mathbf{y},t)\,\mathrm{d}t+\mathbf{\Sigma}(\mathbf{y},t)^{1/2}\,\mathrm{d}\mathbf{B}(t),$ (1a)
$\,\mathrm{d}\mathbf{y}$ $=\mathbf{h}(\mathbf{x},\mathbf{y},t)\,\mathrm{d}t+\mathbf{\Gamma}(\mathbf{y},t)^{1/2}\,\mathrm{d}\mathbf{W}(t).$ (1b)
$\mathbf{f}$ is the drift of the signal process (1a) and $\mathbf{h}$ is the drift of the observation process (1b). The terms $\mathbf{B}(t)\in\mathbb{R}^{n_{\mathbf{x}}}$ and $\mathbf{W}(t)\in\mathbb{R}^{n_{\mathbf{y}}}$ are independent standard Wiener processes. For simplicity, we assume that the signal process noise covariance $\mathbf{\Sigma}(\mathbf{y},t)\in\mathbb{R}^{n_{\mathbf{x}}\times n_{\mathbf{x}}}$ and the observation noise covariance $\mathbf{\Gamma}(\mathbf{y},t)\in\mathbb{R}^{n_{\mathbf{y}}\times n_{\mathbf{y}}}$ are symmetric and positive definite. Throughout the paper, all SDEs, whether written in differential form or in formal derivative form, are understood in the Itô sense [34, 51].
We consider a fixed time horizon $t\in[0,T]$, and aim to characterize the conditional distribution of $\mathbf{x}(t)$ given the available observations. Let $\mathbf{y}_{[0,\tau]}=\{\mathbf{y}(s):0\leq s\leq\tau\}$ denote the observation trajectory up to time $\tau\in[0,T]$. We formulate the filtering and smoothing problems of DA as follows:
  1. 1.
    Filtering: Estimate the conditional distribution of $\mathbf{x}(t)$ given the observational history up to the present $\mathbf{y}_{[0,t]}$. The filtering distribution is denoted by $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,t]})$, with mean $\overline{\mathbf{x}}_{\mathrm{f}}(t)$ and covariance $\mathbf{P}_{\mathrm{f}}(t)$.
  2. 2.
    Smoothing: Estimate the conditional distribution of $\mathbf{x}(t)$ given the full observation trajectory $\mathbf{y}_{[0,T]}$, including the past and the future. The smoothing distribution is denoted by $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,T]})$, with mean $\overline{\mathbf{x}}_{\mathrm{s}}(t)$ and covariance $\mathbf{P}_{\mathrm{s}}(t)$.
It is worth mentioning that the coefficients in (1), including the drifts $\mathbf{f}$ and $\mathbf{h}$, as well as the noise covariances $\mathbf{\Sigma}$ and $\mathbf{\Gamma}$, may depend explicitly on the observation history $\mathbf{y}_{[0,t]}$. Since $\mathbf{y}$ is given in the filter and smoother, its trajectory can be treated as time-varying parameters [52]. This flexibility allows for broad application in which the observed variables actively feed back on and modulate the hidden dynamics; however, for our purposes, we assume the Markovian structure in (1) [51]. For notational simplicity, we may omit the explicit dependence of the model components on the dependent and time variables when the context is clear, and write $\mathbf{f}(\mathbf{x},\mathbf{y},t)$ as $\mathbf{f}$, and similarly for $\mathbf{h}$, $\mathbf{\Sigma}$, and $\mathbf{\Gamma}$.
In general, evolving the full filtering or smoothing distribution is computationally infeasible [52]. Therefore, we focus on estimating the first two moments: the mean and covariance. The idea is that in systems with linear dynamics and Gaussian noise, the conditional distributions are Gaussian and fully described by their mean and covariance, i.e., $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,t]})=\mathcal{N}(\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t))$ and $p(\mathbf{x}(t)\mid\mathbf{y}_{[0,T]})=\mathcal{N}(\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t))$. These exact moments are governed by the classical Kalman–Bucy Filter [53] and continuous-time Rauch–Tung–Striebel (RTS) smoother equations [33].
For general nonlinear systems where closed-form DA solutions are unavailable, we adopt an ensemble-based approach [21, 38]. Here, the conditional distributions are represented by a finite ensemble of size $m$. This can be viewed as a Monte Carlo approximation in which the conditional moments are replaced by empirical ensemble moments. Each ensemble member independently evolves according to (1), but the ensemble constituents are coupled through the empirical mean and covariance.
Let $\mathbf{x}^{(i)}(t)$, $i=1,\dots,m$, denote the ensemble members. The empirical mean $\overline{\mathbf{x}}(t)$ and covariance $\mathbf{P}(t)$ for $t\in[0,T]$ are defined by
$\overline{\mathbf{x}}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{x}^{(i)}(t),\qquad\mathbf{P}(t):=\frac{1}{m-1}\sum_{i=1}^{m}\big(\mathbf{x}^{(i)}(t)-\overline{\mathbf{x}}(t)\big)\big(\mathbf{x}^{(i)}(t)-\overline{\mathbf{x}}(t)\big)^{\top}.$
When needed, we use subscripts $\mathrm{f}$ and $\mathrm{s}$ to distinguish between the filter- and smoother-based ensembles. For brevity, we use $\big(\overline{\mathbf{x}}_{\mathrm{f}},\mathbf{P}_{\mathrm{f}}\big)$ and $\big(\overline{\mathbf{x}}_{\mathrm{s}},\mathbf{P}_{\mathrm{s}}\big)$ to denote the corresponding empirical moments, and do not distinguish them from the true moments unless confusion may arise.
While filtering is the standard approach for online state estimation, many applications, such as Earth system reanalysis [11] and parameter estimation [54, 55], necessitate a retrospective analysis once the full observation trajectory over $[0,T]$, $\mathbf{y}_{[0,T]}$, is available. Based on the filtering quantities $\big(\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\big)$, the smoothing quantities $\big(\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\big)$ incorporate information from future observations $\mathbf{y}_{[t,T]}$, producing more accurate estimates and reducing uncertainty over $[0,T]$. This improvement motivates a systematic forward-backward structure algorithm [21, 56]:
  1. 1.
    Forward pass (filtering): Integrate forward from $t=0$ to $t=T$ to compute the filter ensemble $\mathbf{x}^{(i)}_{\mathrm{f}}(t),i=1,\dots,m$ (forecasting). The filtering mean and covariance $\{\overline{\mathbf{x}}_{\mathrm{f}}(t),\mathbf{P}_{\mathrm{f}}(t)\}_{0\leq t\leq T}$ are obtained.
  2. 2.
    Backward pass (smoothing): Integrate backward-in-time from $t=T$ to $t=0$. Using the filtering estimates as a prior, this pass produces the smoothing ensemble $\mathbf{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$ and therefore its statistics $\{\overline{\mathbf{x}}_{\mathrm{s}}(t),\mathbf{P}_{\mathrm{s}}(t)\}_{0\leq t\leq T}$ (hindcasting).
In our method, we implement this architecture by utilizing the continuous-time ensemble Kalman–Bucy filter (EnKBF) for the forward pass [57], and the EnKBS for the backward pass. The operational details of these methods are presented in the following subsections.
2.2 The Forward Pass: EnKBF
We present a stochastic variant of EnKBF in this subsection, dropping the subscripts $\mathrm{f}$ for notational simplicity. First, assume that the observation drift $\mathbf{h}$ is differentiable with respect to $\mathbf{x}$. Let $\mathbf{h}\approx\mathbf{H}\mathbf{x}$ denote its local linearization, defined through its Jacobian $\mathbf{H}$. The classical discrete-time Kalman filter under the lens of Bayesian inference consists of two consecutive steps [25, 32, 34]: the update or forecast step, which propagates the dynamics to obtain a model-based prior, and the analysis step, which assimilates new observations into the prior, through the likelihood, to produce a posterior via Bayes’ theorem. Translating this viewpoint to the continuous-time setting, we regard the analysis step as incorporating observations via a continuous flow rather than abrupt discrete updates. This is formulated by introducing a fictitious assimilation time interval $s\in[0,1]$ at each physical time $t$ [47, 57, 58]. Specifically, for each ensemble member, we generate a simulated observation increment
$\mathrm{d}\mathbf{y}^{(i)}:=\mathbf{H}\mathbf{x}^{(i)}\,\mathrm{d}t+\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)},$ (2)
where $\{\mathbf{W}^{(i)}\}_{i=1,\ldots,m}$ are mutually independent standard Wiener processes. At each physical time $t$, the analysis step update for the $i$-th ensemble member, $i=1,\dots,m$, is then governed by the fictitious-time SDE
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),\qquad s\in[0,1],$ (3)
with Kalman gain $\mathbf{K}:=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Adding (3) as an additional drift to the “forecast” dynamics (1a) yields the stochastic continuous-time EnKBF formulation with a linear observation operator $\mathbf{H}$ [57, 59]:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right).$ (4)
Treating the simulated observations as random variables corresponding to each ensemble member ensures statistical consistency with the Riccati covariance equation in classical Kalman filter theory; in the absence of the noise term $\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)}$ in (2), the filtering covariance would be systematically underestimated [39, 60].
In practical applications, the observation drift $\mathbf{h}$ is often nonlinear. While we rely on the local linearization $\mathbf{h}\approx\mathbf{H}\mathbf{x}$ to derive (4), we apply the following approximations to obtain a derivative-free framework:
$\mathbf{H}\mathbf{x}^{(i)}\approx\mathbf{h}(\mathbf{x}^{(i)},t),$ (5)
$\mathbf{P}\mathbf{H}^{\top}\approx\mathbf{P}_{\mathbf{x}\mathbf{h}}.$ (6)
Here, $\mathbf{P}_{\mathbf{x}\mathbf{h}}$ denotes the cross-covariance matrix between the ensemble and its observation-operator-evolved counterpart:
$\mathbf{P}_{\mathbf{x}\mathbf{h}}:=\frac{1}{m-1}\sum_{i=1}^{m}\left(\mathbf{x}^{(i)}-\overline{\mathbf{x}}\right)\left(\mathbf{h}(\mathbf{x}^{(i)})-\overline{\mathbf{h}}\right)^{\top},$
where
$\overline{\mathbf{h}}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{h}(\mathbf{x}^{(i)},t).$
Substituting (5)–(6) into (4) yields the practical derivative-free EnKBF formulation for $i=1,\dots,m$:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}_{\mathbf{x}\mathbf{h}}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),$ (7)
where, under the approximation (5), the simulated observation in (2) becomes $\mathrm{d}\mathbf{y}^{(i)}=\mathbf{h}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)}$. We remark that applying (5)–(6) avoids explicit tangent-linear or adjoint computations related to $\mathbf{H}$ [12, 40]. The key assumption behind (6) is
$\mathbf{h}({\mathbf{x}}^{(i)})-\overline{\mathbf{h}}\approx\mathbf{H}\big({\mathbf{x}}^{(i)}-\overline{\mathbf{x}}\big),$
which can be interpreted as a finite-difference approximation of the drift $\mathbf{h}$ when the $i$-th ensemble spread, ${\mathbf{x}}^{(i)}-\overline{\mathbf{x}}$, is sufficiently small.
2.3 The Backward Pass: EnKBS
Given a filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}(t),i=1,\dots,m,$ produced by the forward stochastic EnKBF (7), the smoother integrates backward from $t=T$ to $t=0$ to refine the ensemble statistics over $[0,T]$ in the context of the full observed trajectory $\mathbf{y}_{[0,T]}$. Denote the smoothing ensemble by $\mathbf{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$. Initializing the backward pass at $t=T$ by
$\mathbf{x}_{\mathrm{s}}^{(i)}(T)=\mathbf{x}_{\mathrm{f}}^{(i)}(T),$ (8)
the backward-in-time EnKBS ensemble member dynamics for $i=1,\dots,m$ are given by
$\overleftarrow{\,\mathrm{d}}\!\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\,\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\,\mathrm{d}}\!\mathbf{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}_{\mathrm{f}}^{-1}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\,\mathrm{d}}\!t,$ (9)
where $\mathbf{P}_{\mathrm{f}}$ is the empirical covariance of filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}$, and $\overleftarrow{\,\mathrm{d}}(\cdot)$ denotes the backward differential that defines a backward Itô integral, i.e., $\overleftarrow{\,\mathrm{d}}(\cdot)=-\,\mathrm{d}(\cdot)$ [52, 61]. In (9), we use the same Wiener process paths $\mathbf{B}^{(i)}$ as in the forward filter (7), as justified by the derivations in A.
To clarify the backward-in-time differential $\overleftarrow{\,\mathrm{d}}(\cdot)$ in (9), we state the Euler–Maruyama time discretization that we use in the practical algorithm. Let $\tau>0$ denote the time step and $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\mathbf{x}_{k}=\mathbf{x}(t_{k})$. The forward EnKBF (7) is discretized as
$\mathbf{x}_{\mathrm{f},k+1}^{(i)}=\mathbf{x}_{\mathrm{f},k}^{(i)}+\tau\mathbf{f}\left(\mathbf{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Sigma}^{1/2}\mathbf{B}_{k}^{(i)}+\mathbf{P}_{\mathbf{x}\mathbf{h},k}\mathbf{\Gamma}^{-1}\left(\Delta\mathbf{y}_{k+1}-\Delta\mathbf{y}_{k+1}^{(i)}\right),$ (10)
for $k=0,1,\dots,K-1$, where $\mathbf{B}_{k}^{(i)},\mathbf{W}_{k}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent, $\Delta\mathbf{y}_{k+1}=\mathbf{y}_{k+1}-\mathbf{y}_{k}$ is the observation increment, and
$\Delta\mathbf{y}_{k+1}^{(i)}:=\tau\mathbf{h}\left(\mathbf{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Gamma}^{1/2}\mathbf{W}_{k}^{(i)}$
is the simulated observation increment for the $i$-th member, consistent with the continuous-time definition (2). The backward EnKBS (9) is discretized as
$\mathbf{x}_{\mathrm{s},k}^{(i)}=\mathbf{x}_{\mathrm{s},k+1}^{(i)}-\tau\mathbf{f}\left(\mathbf{x}_{\mathrm{s},k+1}^{(i)},t_{k+1}\right)-\sqrt{\tau}\mathbf{\Sigma}^{1/2}\mathbf{B}_{k}^{(i)}-\tau\mathbf{\Sigma}\mathbf{P}_{\mathrm{f},k+1}^{-1}\left(\mathbf{x}_{\mathrm{s},k+1}^{(i)}-\mathbf{x}_{\mathrm{f},k+1}^{(i)}\right),$ (11)
for $k=K-1,K-2,\dots,0$, with initialization $\mathbf{x}_{\mathrm{s},K}^{(i)}=\mathbf{x}_{\mathrm{f},K}^{(i)}$. Note that the same noise increments $\mathbf{B}_{k}^{(i)}$ from the forward pass are reused in the backward pass.
As a remark, in the EnKBS formulation (9), the backward dynamics do not depend explicitly on the observation increment $\mathrm{d}\mathbf{y}$ as with the EnKBF in (7). The implicit dependence on $\mathbf{y}$ is instead via the filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}(t)$} computed in the forward pass. Under the assumption of independent signal and observation white noises, the information in $\mathbf{y}$ is already encoded in the future state $\mathbf{x}$ and propagates backward to correct the current estimate along the smoother pass. Conditioned on the future $\mathbf{x}$, the Markov property of the system yields that including any observation $\mathbf{y}$ depending on it does not provide additional information in the absence of cross-interacting noise [49, 52, 62].
Detailed derivations of (9) and its formal connection to the discrete-time RTS smoother in the formal limit of a vanishing observation interval are provided in A. In addition, B shows that the proposed EnKBS recovers the exact mean and covariance equation of the continuous-time RTS smoother in the linear-Gaussian setting [33].
2.4 Localization and Inflation
In high-dimensional data assimilation with state dimension $n$, a small ensemble size $m\ll n$ inevitably incurs sampling error and cannot fully resolve all dynamic modes. Specifically, the full covariance matrix is $n\times n$, whereas the empirical ensemble covariance $\mathbf{P}$ is rank-deficient with rank at most m1m-1, which is often insufficient to fully capture the dominant unstable modes of the system [63]. This rank deficiency may cause instability or divergence of the filter and smoother and induce spurious, nonphysical long-range correlations [40, 48]. In many operational applications, affordable ensemble sizes typically range from $O(10)$ to $O(100)$, which can be much smaller than the dimension of the unstable subspace [12]. Increasing $m$ to that scale is often computationally prohibitive.
Two standard regularization techniques, localization and inflation, are widely used to address these issues and improve stability in high-dimensional systems. These ad hoc strategies, to some extent, help explain the practical success of ensemble methods in applications such as numerical weather prediction and the geosciences [42].
Localization
Covariance localization, also called Schur localization, regularizes the empirical covariance by damping long-range state-space correlations and thereby increases its rank [12, 47]. A standard approach applies a Schur (element-wise) product between the empirical covariance $\mathbf{P}$ and a short-ranged correlation matrix $\mathbf{C}_{\mathrm{loc}}\in\mathbb{R}^{n\times n}$:
$\mathbf{P}\leftarrow\mathbf{C}_{\mathrm{loc}}\circ\mathbf{P},$ (12)
where $\circ$ denotes the Schur product of matrices.
By the Schur product theorem [64], $\mathbf{C}_{\mathrm{loc}}\circ\mathbf{P}$ is positive (semi-)definite whenever $\mathbf{C}_{\mathrm{loc}}$ and $\mathbf{P}$ are positive (semi-)definite. A common choice constructs $\mathbf{C}_{\mathrm{loc}}$ from the Gaspari–Cohn fifth-order piecewise rational function $G$ [65],
$G(r)=\begin{cases}1-\dfrac{5}{3}r^{2}+\dfrac{5}{8}r^{3}+\dfrac{1}{2}r^{4}-\dfrac{1}{4}r^{5},&0\leq r<1,\\ 4-5r+\dfrac{5}{3}r^{2}+\dfrac{5}{8}r^{3}-\dfrac{1}{2}r^{4}+\dfrac{1}{12}r^{5}-\dfrac{2}{3r},&1\leq r<2,\\ 0,&r\geq 2,\end{cases}$
for $r\geq 0$. It resembles a Gaussian distribution, while remaining compactly supported and inexpensive to evaluate. Given a localization radius $r_{0}>0$ and a distance $d(i,j)$ between state components $i$ and $j$, we define
$(\mathbf{C}_{\mathrm{loc}})_{i,j}:=G\left(\frac{d(i,j)}{r_{0}}\right),$
so that correlations are truncated beyond $d(i,j)>2r_{0}$.
Accordingly, the localized formulation of the stochastic EnKBF (7) replaces covariance $\mathbf{P}_{\mathbf{x}\mathbf{h}}$ by $\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\mathbf{x}\mathbf{h}}$ [66]:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+(\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\mathbf{x}\mathbf{h}})\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),$ (13)
and, for consistency, we also apply a localization to the backward EnKBS (9):
$\overleftarrow{\,\mathrm{d}}\!\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\,\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\,\mathrm{d}}\!\mathbf{B}^{(i)}+\mathbf{\Sigma}(\mathbf{C}_{\mathrm{loc,2}}\circ\mathbf{P}_{\mathrm{f}})^{-1}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\,\mathrm{d}}\!t.$ (14)
Inflation
With localization, sampling error and spurious correlations are reduced, but still may accumulate over time. A practical solution to this is multiplicative inflation, which rescales ensemble anomalies around the mean by a factor $\delta^{2}>1$ [67]:
$\mathbf{x}^{(i)}\longleftarrow\overline{\mathbf{x}}+\delta\big(\mathbf{x}^{(i)}-\overline{\mathbf{x}}\big),\qquad i=1,\dots,m.$ (15)
The multiplicative inflation (15) aims to compensate for sampling errors that are the indirect consequence of the nonlinearity in the ensemble forecast, which helps to cure the intrinsic error of ensemble filter schemes [12]. In addition, inflation should be applied only for the forward filtering, not the backward smoothing pass. Since sampling errors are already accounted for in the filter, applying inflation in the backward update would result in a suboptimal solution [68].
3 EnKBS for State Estimation, Causal Inference, and Model Identification
In this section, the EnKBS is adopted to study three topics in complex nonlinear dynamical systems. The first study is the state estimation of a relatively high-dimensional chaotic system, which highlights the necessity of regularization and demonstrates the applicability of EnKBS in such systems. The remaining two examples are nonlinear coarse-grained conceptual models, where EnKBS is adopted for advancing causal inference and discovering the hidden model structure and parameters.
In all experiments, we use the Euler–Maruyama scheme for numerical time integration [69]. For the backward smoothing pass, we discretize the formulation (9) and propagate the ensemble backward in time.
3.1 State Estimation in a Relatively High-Dimensional Chaotic System: The Lorenz-96 Model
The first example considers the state estimation of the Lorenz-96 system. Introduced by Edward Lorenz, it was originally designed to mimic the propagation of atmospheric waves on a midlatitude zonal circle [18, 70]. It is now widely used as a benchmark test case in data assimilation. Let $\mathbf{x}=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$. We consider the stochastic Lorenz-96 dynamics:
$\dot{x}_{j}=(x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F+\sigma_{j}\dot{W}_{j},\qquad j=1,\dots,n.$ (16)
Periodic indexing modulo $n$ is applied, i.e., $x_{j+n}=x_{j}$ for all integers $j$. Here, $W_{i}$ are mutually independent Wiener processes and $\dot{x}$ denotes derivative with respect to time $t$.
The quadratic term $(x_{j+1}-x_{j-2})x_{j-1}$ mimics nonlinear advection and preserves the total energy, which is proportional to $\sum_{i}x_{i}^{2}$. The linear term $x_{j}$ represents mechanical or thermal dissipation and provides system damping, and the constant forcing $F$ drives the system into a chaotic regime when it exceeds a certain threshold value. With this parameter choice, the deterministic system has $13$ positive and one neutral Lyapunov exponent, corresponding to the dimension of the unstable-neutral subspace [63, 70].
3.1.1 The Experiment Setup
In this test, we observe every other state component (the even indices) and treat the remaining components as hidden [55]. In other words, the hidden signal process (1a) and observational process (1b) correspond respectively to
$\dot{x}_{j}$ $=(x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F+\sigma_{j}\dot{W}_{j},\qquad j=1,3,5,\dots,n-1,$
$\dot{x}_{j}$ $=(x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F+\sigma_{j}\dot{W}_{j},\qquad j=2,4,6,\dots,n.$
We set the noise variance in the observed processes to $\sigma_{j}^{2}=0.1$, $j=2,4,6,\dots,n$, and the noise variance in the unobserved processes as $\sigma_{j}^{2}=5$, $j=1,3,5,\dots,n-1$. The dimension of the system is $n=40$ and the forcing $F=8$ represents a strongly chaotic regime.
We set the observation rate to be equal to the numerical integration time step $\Delta t=0.005$. A reference trajectory $\mathbf{x}_{\mathrm{ref}}$ is generated over the time window $[0,T]$ with $T=100$, and a segment is shown in Figure 1. We initialize the ensemble by adding small perturbations to the initial reference state. For localization, we use the periodic distance between indices $i$ and $j$:
$d(i,j):=\min\big\{|i-j|,n-|i-j|\big\},\qquad 1\leq i,j\leq n.$
Figure 1: Lorenz-96 reference trajectory over $t\in[75,100]$. The horizontal axis is time, the vertical axis is the state index $j\in\{1,\dots,40\}$, and the color indicates the value of $x_{j}(t)$.
3.1.2 Results
The localization radius varies among $r_{0}\in\{1,2,3,4,5,8,15,18\}$. All figures and tables below use an ensemble size of $m=10\,(\leq 13)$. Tables 1 and 2 report the filter and smoother root-mean-square errors (RMSEs), respectively, as functions of the localization radius $r_{0}$ and the inflation factor $\delta^{2}$. Here, the RMSE is defined by
$\mathrm{RMSE}:=\sqrt{\frac{1}{nK}\sum_{k=1}^{K}\left\|\overline{\mathbf{x}}(t_{k})-\mathbf{x}_{\mathrm{ref}}(t_{k})\right\|^{2}},$
where $K$ is the total number of time steps and $\overline{\mathbf{x}}$ denotes the ensemble mean (filter or smoother). The results show that, with localization and mild inflation, the smoother consistently improves upon the filter on Lorenz-96. While a suitable inflation factor slightly reduces the error, localization has a pronounced impact on the stability and convergence of EnKBS. Compared with EnKBF, EnKBS is more sensitive to ensemble size. Without localization, EnKBS typically requires $m\gtrsim 25$ to remain stable in this setting; for smaller $m$, the backward pass diverges. In contrast, EnKBF remains stable even with ensembles as small as $m=3$. With localization, EnKBS remains stable for much smaller ensembles (down to $m=5$) and still outperforms the corresponding filter estimate (not shown). Figure 2 compares the time trajectories of $x_{1}$ between the reference solution and the EnKBF/EnKBS mean, with parameters $r_{0}=3$ and $\delta^{2}=1.005$.
$\delta^{2}\backslash r_{0}$ 1 2 3 4 5 8 15 18
1 0.714 0.662 0.656 0.660 0.668 0.704 0.734 1.025
1.0001 0.712 0.662 0.656 0.659 0.668 0.704 0.733 1.025
1.001 0.700 0.659 0.654 0.658 0.666 0.703 0.732 1.026
1.005 0.683 0.664 0.654* 0.654 0.661 0.698 0.728 1.033
1.01 NaN 0.762 0.683 0.667 0.671 0.698 0.727 1.041
1.02 NaN NaN NaN NaN NaN 0.848 1.037 1.062
Table 1: Lorenz-96: RMSE of the localized and inflated EnKBF over the evaluation window $[0,100]$. Entries marked as NaN indicate numerical divergence. The symbol * marks the smallest RMSE among the tested parameters.
$\delta^{2}\backslash r_{0}$ 1 2 3 4 5 8 15 18
1 0.583 0.567 0.562 0.574 0.591 0.647 0.686 NaN
1.0001 0.581 0.560 0.561 0.573 0.590 0.646 0.685 NaN
1.001 0.565 0.552 0.555 0.568 0.585 0.643 0.682 NaN
1.005 0.663 0.535 0.531 0.544 0.563 0.630 0.673 NaN
1.01 NaN 0.643 0.528 0.519* 0.541 0.614 0.662 NaN
1.02 NaN NaN NaN NaN NaN 0.766 1.001 NaN
Table 2: Lorenz-96: RMSE of the localized and inflated EnKBS over the evaluation window $[0,100]$. Entries marked as NaN indicate numerical divergence. The symbol * marks the smallest RMSE among the tested parameters.
Figure 2: Trajectories of $x_{1}(t)$ over $t\in[70,100]$ for Lorenz-96. The black curve is the reference solution; the red and green curves are the EnKBF and EnKBS ensemble means, respectively. Shaded bands show the respective $\pm 2$ standard-deviation intervals. Parameters: $m=10$, $r_{0}=3$, and $\delta^{2}=1.005$.
3.2 Causal Inference in a Nonlinear Dyad Model
We incorporate the EnKBS into the recently developed assimilative causal inference (ACI) framework [16]. Unlike traditional approaches based solely on time-averaged statistics or model forecasts, ACI casts causal inference as an inverse problem through Bayesian data assimilation, enabling time-dependent diagnosis of causal dynamics. In this case study, we use a nonlinear dyad model to demonstrate that the EnKBS-ACI framework is able to infer the causal intensity and temporal influence range of a hidden driver from observations of its effect, dynamically identifying the instantaneous cause-and-effect role interplay.
The key idea behind ACI is that, if a variable $\mathbf{y}$ drives another $\mathbf{x}$, then conditioning on the effect $\mathbf{x}$’s future evolution should retrospectively improve the estimate of the driver $\mathbf{y}$’s current state. In ACI, the instantaneous causal influence intensity is quantified by an information-gain metric (ACI metric) based on the relative entropy between the smoother distribution and the real-time filtering distribution. By further evaluating the information gain across lagged smoothers of different assimilating time windows, we can also obtain the (forward) causal influence range (CIR) of a relationship, which characterizes how far the causal signal propagates in time. This CIR is objective, i.e., it does not rely on subjective or empirical threshold choices. Here, using the EnKBS to approximate the filter and smoother distributions removes the need to directly observe the driver for causal inference, as long as its effect is observed.
3.2.1 A Nonlinear Dyad Model with Extreme Events
We consider a nonlinear dyad model that exhibits intermittent extreme events through a trigger–feedback mechanism:
$\dot{u}$ $=(-d_{u}+cv)u+F_{u}+\sigma_{u}\dot{W}_{u},$ (17a)
$\dot{v}$ $=-d_{v}v-cu^{2}+F_{v}+\sigma_{v}\dot{W}_{v}.$ (17b)
Here, $W_{u}$ and $W_{v}$ are independent standard Wiener processes. This physics-constrained low-order model produces extreme events through a trigger–feedback cycle: $v$ modulates the damping of $u$, while $u$ feeds back onto $v$ [71]. Specifically, the growth rate of $u$ is controlled by the time-varying coefficient $d_{u}+cv$. When intermittent growth of $v$ exceeds the threshold $d_{u}/c$, the system enters an anti-damping regime and $u$ grows rapidly. As $u$ amplifies, the term $cu^{2}$ suppresses $v$, which restores damping in $u$ and terminates the burst in its amplitude, ultimately conserving the energy of the system.
3.3 The Experiment Setup
The following parameter values are adopted in (17):
$d_{u}=0.5,\quad F_{u}=1,\quad\sigma_{u}=0.5,\quad c=2,$
$d_{v}=0.5,\quad F_{v}=0.8,\quad\sigma_{v}=1.$
We consider both causal directions, the triggering direction $v\to u$ and the feedback direction $u\to v$. For $v\to u$, the filter and smoother estimate the hidden state $v$ given the observations of $u$. Notice that when $u$ is observed, the $v$ dynamics (17b) become conditionally linear and Gaussian, fitting the conditional Gaussian nonlinear system (CGNS) structure [72, 73]. This provides an analytic filter and smoother that enables a benchmark comparison with the EnKBF and EnKBS [49, 16]. For $u\to v$, we estimate the hidden state $u$ from observations of $v$. Note that this predictability arises through the coupling term $cu^{2}$ in (17b). However, because the map $u\mapsto u^{2}$ is non-injective, conditioning on $v$ alone does not resolve the sign of $u$. As a result, when the true $u$ is close to 0, ensemble members may split into opposite branches, artificially inflating the variance and increasing the estimation error. Hence, we fix a reference solution of length $T=30$ over which $u$ does not change sign for both directions. See the top panel of Figure 3. We observe that inflation does not noticeably change the results in this regime, and we therefore do not apply inflation.
Figure 3: $v\rightarrow u$. CIR and ACI metric computed using the EnKBS with ensemble size $m=50$. Top: true trajectories of $u$ (magenta) and $v$ (blue). The dashed line is the anti-damping threshold $d_{u}/c$. Middle: the CIR length for $v\rightarrow u$. Bottom: the ACI metric as a function of time.
3.3.1 Results
Triggering: $v\rightarrow u$}
Figure 4 compares the RMSE between the EnKBF/EnKBS and CGNS analytic filter/smoother estimates of $v$ as a function of ensemble size $m$. As $m$ increases, the EnKBS ensemble mean and covariance converge to the analytic statistics, with the resulting ACI diagnostics converging accordingly. Figures 3 and 5 show the ACI results with an ensemble size of $m=50$ and $m=10$ respectively. With $m=50$, the EnKBS diagnostics closely match the analytic benchmark. We therefore display only the CIR and ACI metric results with the EnKBS. For $m=10$, the ACI metric does not preserve relative peak magnitudes and underestimates the CIR at the beginning of the rise of $v$, i.e., during the build-up phase of $u$’s extreme event; this is especially pronounced at prolonged or delayed onsets. Nevertheless, it captures the peak locations and yields qualitatively similar patterns for the CIR and ACI metric.
In general, the significant ACI metric coincides with significant $v$ before extreme events of $u$, e.g., $t\approx 7$, $9.5$, and $16$, indicating the strongest instantaneous influence where $v$ acts as the primary anti-damping driver of these extreme events. The CIR peaks at the onset of $v$’s growth, e.g., $t\approx 6.5$, $8$, and $15$, suggesting long-range dependence from $v$ to $u$ during the onset stage of these extreme events, and the triggering conditions of extreme events of $u$ are established well in advance. The CIR then shortens, and the ACI metric drops sharply as the events approach their peaks and begin to decay.
From a data assimilation perspective, the filter cannot detect the triggering conditions at the onset phases unless they are manifested in the observations. In contrast, the smoother is able to clearly improve upon the filter with future information in hindsight. As a result, a larger ACI metric and a longer CIR are obtained. Later, when $u$ begins to strengthen, near-future observations already provide sufficient information for effective smoothing. Once $u$ becomes strong enough, its signal-to-noise ratio enables the filter to effectively capture the relevant triggering dynamics, leaving little room for the smoother to improve estimates. Therefore, the large CIR and ACI metric does not persist in the peak and demise stages of extreme events of $u$.
Figure 4: RMSE comparison between EnKBF/EnKBS and CGNS exact filter/smoother as a function of ensemble size $m$. Red circle: EnKBF. Green square: EnKBS. Red dashed: optimal filter. Green dashed: optimal smoother. The RMSE between the filter/smoother estimates and the truth of $v$ is evaluated over a time window of $[0,500]$.
Figure 5: $v\rightarrow u.$ CIR and ACI metric computed using EnKBS with ensemble size $m=10$. Top: true trajectories of $u$ (magenta) and $v$ (blue). The dashed line is the anti-damping threshold $d_{u}/c$. Middle: the CIR length for $v\rightarrow u$. Bottom: the ACI metric as a function of time.
Feedback: $u\rightarrow v$}
In this regime, experiments use an ensemble size of $m=50$ (based on the results of Figure 4). Figure 6 shows the estimation errors and empirical standard deviations for both the filter and the smoother. During extreme events, the smoother provides a clear correction relative to the filter. No sign-splitting is observed since $u>0$ throughout.
Figure 6: $u\rightarrow v$. EnKBF vs. EnKBS diagnostics with ensemble size $m=50$. Top: true trajectories of $u$ (magenta) and $v$ (blue). Middle: estimation errors of $u$ for the filter (red) and smoother (green), computed as $u_{\mathrm{ref}}-u_{\mathrm{f}}$ and $u_{\mathrm{ref}}-u_{\mathrm{s}}$, respectively. Values closer to 0 (dashed) indicate smaller errors. Bottom: empirical standard deviation of $u$ for the filter (red) and smoother (green).
Figure 7 shows the CIR and ACI metric. Large ACI metric matches the peak of $v$, e.g., $t\approx 7.6$, $9.6$, and $13$. This indicates the strongest causal direction in which $u$ suppresses the growth of $v$. The CIR is relatively long during the onset stages of extreme events of $u$, e.g., $t\approx 6.5$, $8.5$, $10.8$, and $12$, suggesting that, as in the $v\to u$ direction, the $u\to v$ feedback has been established well before extreme events. As $v$ approaches its peak and proceeds towards its demise, the CIR gradually shortens. Changing the ensemble size to $m=100$ or $m=10$ yields quantitatively similar results (not shown).
Figure 7: $u\rightarrow v$. CIR and ACI metric computed using EnKBS with ensemble size $m=50$. Top: true trajectories of $u$ (magenta) and $v$ (blue). The dashed line is the anti-damping threshold $d_{u}/c$. Middle: the CIR length for $u\rightarrow v$. Bottom: the ACI metric as a function of time.
To interpret these patterns, given a time series of $\mathrm{d}v$ in the $v$-equation (17b), the filter infers $u$ through the negative coupling term $cu^{2}$. As a result, the filter cannot anticipate the growth and peak of $u$ unless the subsequent decline of $v$ is observed. In contrast, the smoother can trace back from this decline to correctly estimate $u$. Consequently, both the CIR and the ACI metric become large during the onset stage, and the CIR gradually shortens as the decline of $v$ is imminent. Later, $v$ starts to decay. This phenomenon already provides an informative signal to the filter, and the information flow $u\rightarrow v$ has largely been realized and manifested in the CIR and ACI metric during the onset period. Therefore, the filter can achieve an accurate estimate via dynamical interpolation, so the additional future information becomes negligible. As a result, the ACI metric does not persist after the peak and through the full demise of the extreme events.
3.4 Data-Driven Discovery of Dynamics with EnKBS-Based Sampling
In this subsection, we adapt a causality-based learning algorithm to recover the underlying dynamics of partially observed systems. The key idea is to formulate model learning as an iterative procedure that alternates among conditional sampling, structure identification, and parameter estimation [15]. The procedure consists of three steps:
  1. 1.
    Conditional sampling. Based on partially observed time series, reconstruct trajectories of the hidden variables.
  2. 2.
    Structure identification. Detect which candidate terms are active in the governing equations by quantifying the information transfer from each functional, in a library of functions, to the trajectory of each state.
  3. 3.
    Parameter estimation. Given the identified structure, recover the corresponding coefficients via maximum likelihood estimation.
The EnKBS plays a key role in Step 1. In our approach, the smoother ensemble member trajectories provide conditional samples of the hidden states. We treat these trajectories as synthetic observations to complete the dataset and enable subsequent identification of the model structure and parameters.
We incorporate physics constraints, e.g., energy conservation in quadratic terms, to prevent nonphysical finite-time blow-up and stabilize learning [74, 75]. These constraints are enforced in the parameter-estimation step together with other constraints to facilitate learning convergence.
3.4.1 The Lorenz-84 System
Numerical tests are performed on the Lorenz-84 model [76]:
$\dot{x}$ $=-(y^{2}+z^{2})-a(x-f)+\sigma_{x}\dot{W}_{x},$ (18)
$\dot{y}$ $=-bxz+xy-y+g+\sigma_{y}\dot{W}_{y},$
$\dot{z}$ $=bxy+xz-z+\sigma_{z}\dot{W}_{z}.$
We use the following parameter setting to induce chaotic dynamics:
$a=\frac{1}{4},\qquad b=4,\qquad f=8,\qquad g=1,\qquad\sigma_{x}=\sigma_{y}=\sigma_{z}=0.1.$
Equation (18) is a low-order atmospheric conceptual model: $x$ denotes the intensity of the mid-latitude westerly zonal flow, while $y$ and $z$ represent the cosine and sine longitudinal phases of superimposed large-scale wave patterns. Stochastic forcing is included to represent interactions with unresolved sub-grid processes, while $f$ and $g$ represent symmetric (e.g., by solar heating) and asymmetric (e.g., from topographic influences) cross-latitude heating contrasts, respectively.
3.4.2 The Experiment Setup
We observe $(y,z)$ and treat $x$ as hidden. We use the following 12-term candidate function library:
$1,\;y,\;z,\;y^{2},\;z^{2},\;yz,\;x,\;xy,\;xz,\;xy^{2},\;xz^{2},\;xyz.$
This library matches the CGNS structure, enabling direct comparison with the CGNS optimal sampling benchmark [15, 72]. In the structure-identification step, the constant term is always included by default.
We use the following initial model guess, same as [15]:
$\dot{x}$ $=y^{2}-z^{2}+2+(y^{2}-z^{2})x+\sigma_{x}\dot{W}_{x},$
$\dot{y}$ $=-y-2y^{2}+z^{2}+1+(-y-8z-yz)x+\sigma_{y}\dot{W}_{y},$
$\dot{z}$ $=-z+z^{2}-yz+(8y+z+z^{2})x+\sigma_{z}\dot{W}_{z}.$
The initial guesses for the noise amplitudes of observed variables are set to $\sigma_{y}=\sigma_{z}=1$, while the noise of the hidden $x$ is treated as known and fixed at $\sigma_{x}=0.1$ for simplicity. The time window length is $T=500$ with numerical time step $\Delta t=0.001$. We run $120$ learning iterations. For EnKBS-based sampling, we use an ensemble size of $m=50$. Using $m=100$ or $m=10$ produces similar learning behavior, except that the $m=10$ case typically converges a few iterations later.
3.4.3 Results
Figure 8 shows that, with EnKBS-based sampling in Step 1, the iterative procedure identifies the correct model structure, converges, and recovers the hidden trajectory $x(t)$. In particular, Figure 8(a) shows that the sampled trajectory of the unobserved variable $x$ gradually approaches the truth as iterations proceed.
Figure 8: Learning progress for Lorenz-84 using EnKBS-based sampling. (a) Sampled trajectories of the hidden state $x$ at selected iterations compared with the truth; iteration 0 corresponds to the initial guess. (b) Structure error measured by Frobenius norm $\|\mathbf{C}-\mathbf{C}_{\mathrm{stable}}\|_{F}$. (c) Evolution of the estimated parameter $\hat{b}$ toward the true value $b$.
To quantify structure convergence, we introduce a causation entropy indicator matrix $\mathbf{C}\in\{0,1\}^{N\times M}$ with system dimension $N=3$ and the number of candidate functions $M$ [15]. Each entry is obtained by thresholding the causation entropy:
$\mathbf{C}_{i,j}=\begin{cases}1,&\mathrm{CE}_{i,j}\geq r,\\ 0,&\mathrm{CE}_{i,j}<r,\end{cases}\qquad r=10^{-3},$
where $\mathrm{CE}_{i,j}$ denotes the causation entropy from the $j$-th candidate function to the $i$-th state equation [77, 78]. We then compute the Frobenius norm of the matrix difference $\|\mathbf{C}-\mathbf{C}_{\mathrm{stable}}\|_{F}$, where $\mathbf{C}_{\mathrm{stable}}$ denotes the indicator matrix after the identified structure has stabilized to the true model structure (18). The decay in Figure 8(b) indicates that the model structure is fully identified after about 10 iterations.
Figure 8(c) shows that the estimate of $\hat{b}$ converges to the true value $b$ after roughly 60 iterations. After that point, the remaining coefficients show no noticeable changes. The final parameter estimates at the 120th iteration are listed in Table 3. The identified coefficients are close to the true values.
$\theta^{x}_{x}$ $\theta^{y}_{y}$ $\theta^{z}_{z}$ $\theta^{x}_{y^{2}}$ $\theta^{x}_{z^{2}}$ $\theta^{y}_{xz}$
Truth -0.25 -1 -1 -1 -1 -4
Identified -0.2459 -1.0125 -1.0037 -1.0023 -1.0011 -3.9732
$\theta^{y}_{xy}$ $\theta^{z}_{xy}$ $\theta^{z}_{xz}$ $\theta^{x}_{1}$ $\theta^{y}_{1}$ $\theta^{z}_{1}$
Truth 1 4 1 2 1 0
Identified 1.0023 3.9732 1.0011 1.9946 1.0088 -0.0061
Table 3: Coefficients in the true Lorenz-84 system (18) and the identified model at the 120th iteration. The identified model has the same structure as the truth. Here, $\theta^{x}_{xy}$ denotes the coefficient of the feature xyxy in the $x$-equation, and analogous notation is used for other terms.
Figure 9 compares trajectories and two key statistics generated by the true system and the identified model. When generating the time series, we use the same random seed and initial conditions for the true system and the identified model. The two trajectories match closely up to about $t\approx 10.5$, demonstrating short-term path-wise consistency. Due to chaos, we do not expect long-term point-wise agreement over the full window. Moreover, the identified model reproduces both the probability density function (PDF) and the temporal autocorrelation function (ACF), indicating statistical agreement with the truth.
Figure 9: EnKBS-based sampling. Trajectory and statistics comparison between the true Lorenz-84 system (blue) and the identified model (red). In addition to short-term path-wise agreement, the identified model reproduces the PDF and the temporal ACF of the truth.
Finally, Figures 10 and 11 show the benchmark results using CGNS optimal sampling. The overall learning behavior is similar to the EnKBS-based results in Figures 8 and 9.
Figure 10: Learning progress for Lorenz-84 using CGNS optimal sampling (benchmark). (a) Sampled trajectories of the hidden state $x$ at selected iterations compared with the truth; iteration 0 corresponds to the initial guess. (b) Structure error measured by Frobenius norm $\|\mathbf{C}-\mathbf{C}_{\mathrm{stable}}\|_{F}$. (c) Evolution of the estimated parameter $\hat{b}$ toward the true value $b$.
Figure 11: CGNS optimal sampling. Trajectory and statistics comparison between the true Lorenz-84 system (blue) and the identified model (red). In addition to short-term path-wise agreement, the identified model reproduces the PDF and the temporal ACF of the truth.
4 Conclusions and Discussion
In this paper, we formulate a systematic ensemble Kalman–Bucy smoother (EnKBS) framework for state estimation in nonlinear dynamical systems. The proposed method consists of a forward–backward architecture: a forward filtering pass for online estimation, followed by a backward smoothing pass for retrospective refinement. To handle general nonlinear systems, we adopt an ensemble-based approximation in which conditional moments are replaced by empirical ensemble moments, avoiding tangent-linear and adjoint models. The forward pass assimilates observations to produce filtering estimates, while the backward pass treats the filtered trajectory as a prior and incorporates future information to refine the state estimates.
Through numerical experiments, we show that the EnKBS consistently improves the filtered estimates by reducing error and uncertainty, converging asymptotically to the optimal smoother when an analytic benchmark is available. We further demonstrate that covariance localization and inflation are essential for numerical stability in high-dimensional settings. With regularization, spurious long-range correlations and rank deficiency are mitigated, preventing numerical divergence in high dimensions. Beyond state estimation, we integrate EnKBS into a Bayesian causal inference framework and a causality-based iterative learning algorithm for partially observed systems [16, 15]. With an ensemble size of $O(10)$, EnKBS provides effective state estimates and trajectory samples of unobserved variables in nonlinear systems without analytic smoothers, enabling causality diagnostics around extreme events and supporting model discovery.
As a final remark, although measurements are typically available at discrete times, the continuous-time EnKBS provides a complementary perspective to discrete-time and mixed-time smoothers [53, 57]. This viewpoint is consistent with the continuous-time nature of the underlying physics, and is particularly natural when the goal is to discover dynamical mechanisms. We interpret observed variables as components of the same continuous-time system, rather than as external analysis updates. Assimilation then exploits intrinsic coupling and information transfer within the system, using observed trajectories to constrain the inferred hidden states.
Future work is twofold: expanding the EnKBS methodology and its dynamical discovery applications. On the methodology side, while EnKBS performs well on test systems in this paper, these examples largely involve quadratic or conditionally quadratic structures. A natural next step is to evaluate EnKBS on more realistic, high-dimensional, and strongly nonlinear models. In particular, we can study accuracy and convergence as well as filter/smoother degeneracy together with possible mitigation strategies. On the application side, further experiments and analyses are needed to clarify the mechanisms behind causal information flow and causal reversals in driver-feedback cycles. For example, alongside the forward CIR, we can investigate with the EnKBS the backward CIR to assess the causal precursors that influenced the currently observed event, enabling attribution of significant events in general complex systems [31].
Acknowledgment
The research of N.C. is funded by the Army Research Office W911NF-23-1-0118. Z.J and M.A. are partially supported as research assistants under this grant. M.A. is also supported in part by NSF Award DMS-2023239. The work of S.R. has been partially funded by Deutsche Forschungsgemeinschaft (DFG) – Project-ID 318763901 – SFB1294.
Appendix A Formal Connection to Discrete-time Formulations
In this appendix, we show how the continuous-time stochastic EnKBF formulation (7) and the EnKBS (9) connect to the standard discrete-time Kalman filter and the RTS smoother [32, 33]. By discretizing the continuous-time equations (1), we show that the continuous-time formulations are consistent with the discrete-time filter and smoother up to $O(\tau)$, where $\tau$ is the time increment between observations (assumed uniform, without loss of generality). Starting from the discrete-time filter and smoother, formally taking the limit $\tau\to 0$ yields the continuous-time formulations. As a nonrigorous argument, this is sufficient to highlight the equation structure while avoiding technical details.
Fix a small observation time step $\tau>0$ and define $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\mathbf{x}_{k}=\mathbf{x}(t_{k})$. We consider a first-order semi-implicit Euler–Maruyama type discretization of the nonlinear continuous model (1) [20]:
$\mathbf{x}_{k+1}$ $=\mathbf{x}_{k}+\tau\mathbf{f}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\mathbf{B}_{k},$ (19a)
$\Delta\mathbf{y}_{k+1}$ $=\tau\mathbf{h}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\mathbf{W}_{k+1},$ (19b)
where $\mathbf{B}_{k}$ and $\mathbf{W}_{k+1}$ are independent standard Gaussian vectors, and $\Delta\mathbf{y}_{k+1}=\mathbf{y}_{k+1}-\mathbf{y}_{k}$.
Under standard regularity assumptions, we linearize the drifts as
$\mathbf{f}_{k}=\mathbf{F}_{k}\mathbf{x}_{k}+O(\tau),\qquad\mathbf{h}_{k+1}=\mathbf{H}_{k+1}\mathbf{x}_{k+1}+O(\tau).$ (20)
Plugging (20) into (19), we obtain
$\mathbf{x}_{k+1}$ $=(\mathbf{I}+\tau\mathbf{F}_{k})\mathbf{x}_{k}+\sqrt{\tau}\mathbf{\Sigma}_{k}^{1/2}\mathbf{B}_{k}+O(\tau^{2}),$ (21a)
$\Delta\mathbf{y}_{k+1}$ $=\tau\mathbf{H}_{k+1}\mathbf{x}_{k+1}+\sqrt{\tau}\mathbf{\Gamma}_{k+1}^{1/2}\mathbf{W}_{k+1}+O(\tau^{2}).$ (21b)
We use the notation $(\cdot)_{k\mid\ell}$ for quantities at time $t_{k}$ conditioned on observations up to time $t_{\ell}$. In this notation, $\mathbf{x}_{k+1\mid k}$ is the prior obtained by dynamical forecast from $t_{k}$ to $t_{k+1}$, and $\mathbf{x}_{k+1\mid k+1}$ is the posterior after assimilating the observation at $t_{k+1}$. Ignoring higher order terms $O(\tau^{2})$, (21) leads to the following ensemble Kalman filter analysis update for $i=1,\dots,m$ [39, 59]:
$\mathbf{x}_{k+1\mid k+1}^{(i)}=\mathbf{x}_{k+1\mid k}^{(i)}+\mathbf{K}_{k+1}\left(\Delta\mathbf{y}_{k+1}-\tau\mathbf{H}_{k+1}\mathbf{x}_{k+1\mid k}^{(i)}-\sqrt{\tau}\mathbf{\Gamma}^{1/2}_{{k+1}}\mathbf{W}^{(i)}_{{k+1}}\right)$, (22)
with $\sqrt{\tau}\mathbf{W}^{(i)}_{{k+1}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ a discrete Brownian increment and $\mathbf{K}_{k+1}$ the ensemble-based Kalman gain:
$\mathbf{K}_{k+1}:=\mathbf{P}_{k+1\mid k}\left(\mathbf{H}_{k+1}\right)^{\top}\left(\tau\mathbf{H}_{k+1}\mathbf{P}_{k+1\mid k}(\mathbf{H}_{k+1})^{\top}+\mathbf{\Gamma}_{{k+1}}\right)^{-1}.$
Here, $\mathbf{P}_{k+1\mid k}$ is the empirical covariance of the ensemble prior $\mathbf{x}_{k+1|k}^{(i)}$. From the analysis step of the Kalman filter, we have [12, 32]
$\mathbf{P}_{k+1|k+1}=\mathbf{P}_{k+1\mid k}-\tau\mathbf{K}_{k+1}\mathbf{H}_{k+1}\mathbf{P}_{k+1\mid k}=\mathbf{P}_{k+1\mid k}+O(\tau).$
Together with the following asymptotic relations:
$\mathbf{K}_{k+1}=\mathbf{P}_{k+1\mid k+1}\mathbf{H}_{k+1}^{\top}\mathbf{\Gamma}_{k+1}^{-1}+O(\tau),\qquad\frac{\mathbf{x}_{k+1\mid k+1}^{(i)}-\mathbf{x}_{k+1\mid k}^{(i)}}{\tau}\rightarrow\frac{\,\mathrm{d}\mathbf{x}_{t}^{(i)}}{\,\mathrm{d}s},$
formally taking $\tau\to 0$ in (22) yields the fictitious-time assimilation equation (3). Hence, the stochastic EnKBF formulation (7) for filter-analyzed ensemble members follows by now integrating the derivative-free formulation of (3) (see (5)–(6)) in the forecast dynamics (1a) as an additional drift term [57].
For the smoother, we start from the discrete-time fixed-interval ensemble RTS smoother. Neglecting higher order terms $O(\tau^{2})$, the ensemble-based RTS smoother for (21) reads as [26]
$\mathbf{x}_{k\mid K}^{(i)}=\mathbf{x}_{k\mid k}^{(i)}+\mathbf{C}_{k}\Big(\mathbf{x}_{k+1\mid K}^{(i)}-\mathbf{x}_{k+1\mid k}^{(i)}\Big),\qquad i=1,\dots,m,$ (23)
where the smoothing gain is given by
$\mathbf{C}_{k}:=\mathbf{P}_{k|k}(I+\tau\mathbf{F}_{k}^{\top})\big(\mathbf{P}_{k+1\mid k}\big)^{-1},$ (24)
and the Kalman filter prior covariance by [12, 32]
$\mathbf{P}_{k+1\mid k}=\mathbf{P}_{k\mid k}+\tau\left(\mathbf{F}_{k}\mathbf{P}_{k\mid k}+\mathbf{P}_{k\mid k}\mathbf{F}_{k}^{\top}+\mathbf{\Sigma}_{k+1}\right).$
Hence, its inverse admits the following Neumann series expansion for sufficiently small $\tau$ [64, 79]:
$\big(\mathbf{P}_{k+1\mid k})^{-1}=(\mathbf{P}_{k\mid k})^{-1}-\tau(\mathbf{P}_{k\mid k})^{-1}\left(\mathbf{F}_{k}\mathbf{P}_{k\mid k}+\mathbf{P}_{k\mid k}\mathbf{F}_{k}^{\top}+\mathbf{\Sigma}_{k+1}\right)(\mathbf{P}_{k\mid k})^{-1}+O(\tau^{2}).$ (25)
Plugging (25) into (24) and expanding again, we obtain
$(\mathbf{C}_{k})^{-1}=\mathbf{I}+\tau\big(\mathbf{F}_{k}+\mathbf{\Sigma}_{k+1}(\mathbf{P}_{k\mid k})^{-1}\big)+O(\tau^{2}).$ (26)
Denote the smoother correction for the $i$-th ensemble member by $\mathbf{e}_{k}^{(i)}:=\mathbf{x}_{k|K}^{(i)}-\mathbf{x}_{k|k}^{(i)}$. Rearranging (23) gives
$\mathbf{e}_{k+1}^{(i)}=(\mathbf{C}_{k})^{-1}\mathbf{e}_{k}^{(i)}-\left(\mathbf{x}^{(i)}_{k+1\mid k+1}-\mathbf{x}^{(i)}_{k+1\mid k}\right)$. (27)
At the formal limit $\tau\rightarrow 0$, the filter analysis increment at time $t_{k+1}$ satisfies
$\mathbf{x}^{(i)}_{k+1\mid k+1}-\mathbf{x}^{(i)}_{k+1\mid k}\rightarrow\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)},$
where $\mathrm{d}\mathbf{I}^{(i)}:=\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}$ denotes the innovation process, with simulated observation $\mathrm{d}\mathbf{y}^{(i)}$ from (2). Omitting now the subscripts $(\cdot)_{k+1}$ or $(\cdot)_{k+1|k+1}$ in the continuous-time notation, taking the formal limit $\tau\to 0$ and combining (26) and (27) yields [20]
$\mathrm{d}\mathbf{e}^{(i)}=\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}\big)\mathbf{e}^{(i)}\,\mathrm{d}t-\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)}.$ (28)
Here, $\mathbf{e}^{(i)}=\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}$. Similar to the EnKBF, this can also be regarded as an assimilation SDE in a fictitious time, with information from the future state estimates of $\mathbf{x}$ rather than observations from $\mathbf{y}$. Here, the term $\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\,\mathrm{d}\mathbf{I}^{(i)}$ counteracts the correction in the forward filter pass. Replacing then the tangent linear drift term $\mathbf{F}\mathbf{x}_{\mathrm{s}}^{(i)}$ with the full dynamics $\mathbf{f}(\mathbf{x}_{\mathrm{s}}^{(i)},t)$ and adding (28) to the filter member equation (7) finally yields the EnKBS dynamics (9):
$\mathrm{d}\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\,\mathrm{d}t.$ (29)
Here, $\mathbf{P}=\mathbf{P}_{\mathrm{f}}$ is the empirical filtering covariance in the continuous-time notation ($\mathbf{P}_{k\mid k}\to\mathbf{P}_{\mathrm{f}},\tau\to 0$). Note that since the smoother dynamics $\mathrm{d}\mathbf{x}_{\mathrm{s}}^{(i)}$ are obtained by directly adding the correction $\mathrm{d}\mathbf{e}^{(i)}$ to the forward filter dynamics $\mathrm{d}\mathbf{x}_{\mathrm{f}}^{(i)}$, the same Wiener process path $\mathbf{B}^{(i)}$ from the filter is naturally preserved in the smoother dynamics 29.
Appendix B Moment Consistency of the EnKBS in the Linear-Gaussian Regime
In this appendix, we demonstrate that the EnKBS equation (9) recovers the classical continuous-time RTS smoother mean and covariance equations for linear-Gaussian systems [33, 56]. We note that all empirical covariance matrices are stochastic objects that depend on the observation path $\mathbf{y}_{[0,T]}$ and the driving noises $\mathbf{B}^{(i)}$, $\mathbf{W}^{(i)}$ for finite ensemble sizes. The analysis below is therefore carried out in a mean-field limit $m\to\infty$, where empirical moments are identical to the true conditional moments [80].
Consider the continuous-time linear-Gaussian setting where the drift and observation maps in (1) are linear: $\mathbf{f}=\mathbf{F}\mathbf{x}$ and $\mathbf{h}=\mathbf{H}\mathbf{x}$. Let $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}:=\mathbf{x}_{\mathrm{s}}^{(i)}-\overline{\mathbf{x}}_{\mathrm{s}}$ and $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}:=\mathbf{x}_{\mathrm{f}}^{(i)}-\overline{\mathbf{x}}_{\mathrm{f}}$ denote the smoother and filter anomalies, respectively. By averaging over ensemble members $i$ in the EnKBS member equation (9) and taking a mean-field (infinite ensemble) limit $m\rightarrow\infty$ [80], we recover the exact mean equation of the RTS smoother:
$\mathrm{d}\overline{\mathbf{x}}_{\mathrm{s}}=\mathbf{F}(t)\overline{\mathbf{x}}_{\mathrm{s}}\,\mathrm{d}t+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big(\overline{\mathbf{x}}_{\mathrm{s}}-\overline{\mathbf{x}}_{\mathrm{f}}\big)\,\mathrm{d}t.$ (30)
Subtracting (30) from the EnKBS member equation (9) yields the anomaly dynamics:
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}=\left((\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}})\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\right)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}.$ (31)
We now define the smoother posterior covariance as $\mathbf{P}_{\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$ and the filter–smoother cross-covariance as $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$. Applying Itô’s product rule to $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$ yields [51, 50]:
$\mathrm{d}\big(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big)=(\,\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}(\,\mathrm{d}\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}+\mathbf{\Sigma}\,\mathrm{d}t,$ (32)
where $\mathbf{\Sigma}\,\mathrm{d}t=\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})^{\top}\big]$ emerges as the Itô correction term. Taking the expectation gives the continuous-time covariance equation:
$\dot{\mathbf{P}}_{\mathrm{s}}=\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)\mathbf{P}_{\mathrm{s}}+\mathbf{P}_{\mathrm{s}}\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)^{\top}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{P}_{\mathrm{f},\mathrm{s}}-\mathbf{P}_{\mathrm{f},\mathrm{s}}^{\top}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\Sigma}+\mathbf{\Sigma}.$ (33)
Therefore, the standard continuous-time RTS covariance equation (33) is recovered if
$\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}.$ (34)
Rearranging (34) gives $\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\hat{x}}_{\mathrm{f}}^{(i)})^{\top}\big]=\mathbf{0}$, meaning RTS smoother covariance consistency is achieved if the smoother decreases the ensemble posterior uncertainty beyond the filter solution along directions where the smoother anomaly gain, $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}$, is uncorrelated or orthogonal to the filter anomaly $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}$.
We now show that (34) holds when the stochastic EnKBF formulation is adopted. For the stochastic EnKBF (7), the filter anomalies evolve according to
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}=(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}-\mathbf{K}\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)},$ (35)
where $\mathbf{K}=\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Recall that both the filter anomaly (35) and the smoother anomaly (31) are driven by the exact same system noise $\mathrm{d}\mathbf{B}^{(i)}$. We apply Itô’s product rule to $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$, where the independent observation noise $\mathrm{d}\mathbf{W}^{(i)}$ is uncorrelated with $\mathrm{d}\mathbf{B}^{(i)}$ and the shared system noise generates a non-zero Itô correction term $\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})^{\top}\big]=\mathbf{\Sigma}\,\mathrm{d}t$. Taking then an expectation, we obtain the evolution equation for the cross-covariance $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$:
$\dot{\mathbf{P}}_{\mathrm{f},\mathrm{s}}$ $=\mathbb{E}\left[(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\right]+\mathbb{E}\left[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\Big(\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\Big)^{\top}\right]+\mathbf{\Sigma}$
$=(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{P}_{\mathrm{f},\mathrm{s}}+\mathbf{P}_{\mathrm{f},\mathrm{s}}\big(\mathbf{F}^{\top}+\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\Sigma}\big).$ (36)
On the other hand, the standard continuous-time Riccati equation for the filter covariance $\mathbf{P}_{\mathrm{f}}$ can be written as [53, 34]
$\dot{\mathbf{P}}_{\mathrm{f}}=(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{P}_{\mathrm{f}}+\mathbf{P}_{\mathrm{f}}\big(\mathbf{F}^{\top}+\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\Sigma}\big).$ (37)
Comparing (36) and (37), both $\mathbf{P}_{\mathrm{f},\mathrm{s}}$ and $\mathbf{P}_{\mathrm{f}}$ satisfy the ODE:
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}(\cdot)=(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{(}\cdot)+(\cdot)\big(\mathbf{F}^{\top}+\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\Sigma}\big).$ (38)
Since the smoother ensemble is “initialized” identically to the filter one at the terminal time $t=T$, we have $\mathbf{P}_{\mathrm{f},\mathrm{s}}(T)=\mathbf{P}_{\mathrm{f}}(T)$. Therefore, under appropriate regularity conditions, the solution of (38) is unique and $\mathbf{P}_{\mathrm{f},\mathrm{s}}\equiv\mathbf{P}_{\mathrm{f}}$ for all $t$.
Substituting $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ back into (33), the smoother covariance equation simplifies to
$\dot{\mathbf{P}}_{\mathrm{s}}=\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)\mathbf{P}_{\mathrm{s}}+\mathbf{P}_{\mathrm{s}}\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)^{\top}-\mathbf{\Sigma}.$ (39)
This is the exact continuous-time RTS smoother covariance equation, validating the moment consistency of the EnKBS framework.
We emphasize that recovering the RTS smoother covariance equation relies on the stochastic EnKBF formulation (9). For other deterministic EnKBF formulations based on ensemble transformations [81, 82], $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ does not generally hold. This theoretical mismatch is caused by the coupling with the filter member $\mathbf{x}_{\mathrm{f}}^{(i)}$ in the EnKBS formulation (9). For example, consider a continuous-time deterministic EnKBF formulation [60, 57]:
$\mathrm{d}\mathbf{x}^{(i)}_{\mathrm{f}}=\mathbf{F}\mathbf{x}^{(i)}_{\mathrm{f}}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\frac{\mathbf{H}\mathbf{x}^{(i)}_{\mathrm{f}}+\mathbf{H}\overline{\mathbf{x}}_{\mathrm{f}}}{2}\,\mathrm{d}t\right).$ (40)
Its anomalies obey the dynamics with a half-gain drift (cf. (35)):
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}=\left(\mathbf{F}-\frac{1}{2}\mathbf{K}\mathbf{H}\right)\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}.$ (41)
Applying Itô’s product rule and taking the expectation, we analogously obtain
$\dot{\mathbf{P}}_{\mathrm{f},\mathrm{s}}=\Big(\mathbf{F}-\frac{1}{2}\mathbf{K}\mathbf{H}\Big)\mathbf{P}_{\mathrm{f},\mathrm{s}}+\mathbf{P}_{\mathrm{f},\mathrm{s}}\Big(\mathbf{F}^{\top}+\mathbf{P}_{\mathrm{f}}^{-1}\mathbf{\Sigma}\Big).$ (42)
Clearly, (42) does not match the continuous-time Riccati equation (37) for the filter covariance $\mathbf{P}_{\mathrm{f}}$. To see the discrepancy explicitly, we define the difference matrix $\mathbf{D}:=\mathbf{P}_{\mathrm{f},\mathrm{s}}-\mathbf{P}_{\mathrm{f}}$. Subtracting (37) from (42) yields
$\dot{\mathbf{D}}=\Big(\mathbf{F}-\frac{1}{2}\mathbf{K}\mathbf{H}\Big)\mathbf{D}+\mathbf{D}\Big(\mathbf{F}^{\top}+\mathbf{P}_{\mathrm{f}}^{-1}\mathbf{\Sigma}\Big)+\frac{1}{2}\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathbf{H}\mathbf{P}_{\mathrm{f}}.$ (43)
Although $\mathbf{D}(T)=\mathbf{0}$, the positive semi-definite term $\frac{1}{2}\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathbf{H}\mathbf{P}_{\mathrm{f}}$ generally forces $\mathbf{D}(t)\neq\mathbf{0}$ for $t<T$. Hence, $\mathbf{P}_{\mathrm{f},\mathrm{s}}\neq\mathbf{P}_{\mathrm{f}}$ for this deterministic EnKBF. Consequently, the resulting smoother ensemble covariance $\mathbf{P}_{\mathrm{s}}$ will deviate from the RTS smoother covariance equation (39).
Nevertheless, despite this theoretical mismatch, we have tested this deterministic EnKBF (40) as a forward pass in our framework. It yields results similar to those of the stochastic EnKBF (7) across all numerical experiments presented in this paper (not shown). In addition, for either choice of the forward EnKBF, the smoother consistently improves upon the filter by reducing both the estimation error and the ensemble spread (with behavior similar to that shown in Figure 6).